Given that in an economy, , I, MS =300, Mt = 0.4Y, and Mz=125-200r where, Y= income, C= consumption, I= investment, MS= money supply, Mt= transactional-precautionary money demand, Mz= speculative money demand and r= interest rate. Calculate;
The equilibrium level of income and interest rate in this economy. (2 marks)
The level of C, I, Mt, and Mz when the economy is in equilibrium
1
Expert's answer
2021-07-29T11:29:02-0400

The complete question is:

Given that in an economy, Given that in an economy, C = 102+0.7Y, I=150-100r, MS =300, Mt = 0.4Y, and Mz=125-200r where, Y= income, C= consumption, I= investment, MS= money supply, Mt= transactional-precautionary money demand, Mz= speculative money demand and r= interest rate. Calculate;

1. The equilibrium level of income and interest rate in this economy.

2. The level of C, I, Mt, and Mz when the economy is in equilibrium.00r


Solution

1)

"Y=C+I\\Y=102+0.7Y+150-100r\\Y-0.7Y=252-100r\\0.3Y=252-100r\\Y=840-333.3r\\"


"MS=Mt+Mz\\300=0.4Y+125-200\\0.4Y=300-125+200r\\0.4Y=175+200r\\Y=437.5+500r"


"437.5+500r=840-333.33r\\833.3r=402.5\\r=0.48\\Y=840-333.3(0.48)\\=840-160\\=680"


2)

"C=102+0.7(680)\\=102+476\\=578\\I=150-100(0.48)\\=150-48\\=102\\Mt=0.4(680)\\=272\\Mz=125-200(0.48)\\=125-96\\=29"